OpenStudy (anonymous):

x= 18x+28 _________ 2(x+14) What of the following is a solution for x? A.) x=7 B.)x=-2 C.)x=-14 D.)x=2

OpenStudy (anonymous):

x(2x+28) = 18x +28 2x^2+28x = 18x + 28 2x^2 +28x-18x -28 = 0 2x^2 +10x -28 =0 x = -7 or x =2

OpenStudy (anonymous):

So its C

OpenStudy (anonymous):

no, its D

OpenStudy (anonymous):

x can be either -7 or 2 not (-7)*(2)

OpenStudy (anonymous):

ohh okay thanks

OpenStudy (anonymous):

np